A multiple choice quiz has 400 questions, each question with five possible answers of which one is correct answer. What is the probability that a sheer guess yields exactly 80 correct answers using normal approximation.  


1
Expert's answer
2021-05-27T11:55:28-0400

"X~"~ bin(n, p)

where,

n= 400

p= 1/5 = 0.2

q = 1- p = 0.8


We can say that the sampling distribution of X can be approximately normal with :

mean = "np=(400\times 0.2)=80"


Standard deviation "sd=\sqrt{npq}=\sqrt{400\times 0.2\times 0.8}=8"


Now, the probability that a guesswork yields exactly 80 correct answers is :

"=P(X=80)\\=P(79.5\leq X\leq 80.5)"

"\approx 0.0498" [TI84 function : normalcdf (79.5,80.5,80,8) ]